### v1

@@ 3.2.1 Sample size allocation based on information quantity
$A_{ij}=\exp(-\gamma||f_{i}-f_{j}||^{2}),i,j=1,\cdots,n,$ (1)
where $\gamma$ is a parameter of the Gaussian kernel function, and $f_{i}$ is the feature of the $i$th sample in $D$. The rank of $A$ represents the maximum number of linearly independent samples in $D$ and can reflect the magnitude of the information content within the training set. We use singular value decomposition (SVD) to estimate the effective rank of $A$ as follows:
@@
3.2.2 Sample selection based on spatial distribution
@@
Sample selection. Eq. (6) prevents the selected samples from being overly concentrated in both Euclidean and feature spaces, but it’s an NP-hard problem. Hence we use efficient greedy algorithm to solve Eq. (6), i.e. select the most representative $k$ samples from the training set $D$. In each greedy searching step, we randomly select samples (the set is denoted as $R$) from $D\backslash M$ to speed up the searching process. The definition of $R$ is:
@@ 22      $M_{1:t}=M_{1:t-1}\cup\{M_{t}\}$;
3.3 CL framework based on prompt learning
@@
Prompt module. We design the prompt module with reference to QFormer (Li et al., 2023). It consists of two MLPs (denoted as $\mathrm{MLP_{in}}:R^{3}\rightarrow R^{d}$ and $\mathrm{MLP_{out}}:R^{d}\rightarrow R^{d^{\prime}}$), $k_{q}$ learnable prompts (denoted as $Q\in R^{k_{q}\times d}$), and $N$ attention modules (denoted as $\mathrm{Attn}:R^{d}\rightarrow R^{d}$), as shown in Fig. 3 (c). Here, $\mathrm{MLP_{in}}$ maps the raw point cloud (denoted as $x\in R^{N_{x}\times 3}$) into a high-dimensional space, which then serves as the input to the attention modules. The learnable prompts $Q$ store domain-specific information from each training set and act as the Query in the attention modules. The attention modules $\mathrm{Attn}$ facilitate interaction between $\mathrm{MLP_{in}}(x)$ and $Q$ to extract domain-specific information related to $x$. $\mathrm{MLP_{out}}$ adjusts the dimensionality of the attention module outputs, making it easier to embed prompt information into the backbone network. The dimension $d^{\prime}$ is tailored to accommodate different backbone networks. The extraction of domain-specific information from raw point cloud $x$ by the prompt module is formally expressed as:
@@
Determining how to integrate learnable prompts into the backbone network is a critical problem. Existing prompt learning methods are primarily designed for transformer-based networks. However, most deep learning-based PCPR methods rely on sparse convolution or PointNet-style architectures, requiring the insertion location of prompts to be tailored to the specific structure of the backbone network. For Minkloc3D, a PCPR method based on sparse convolution, this study directly serializes $\mathrm{P}(x)$ with the raw point cloud and uses it as the network input. For PCPR models based on PointNet, such as PointNetVLAD and PatchAugNet, $\mathrm{P}(x)$ is directly added to the intermediate features of the backbone network. Crucially, our method requires no modifications to backbone networks. The insertion locations of the prompt module are visualized in Fig. 3.3.
Fig. 4.: Location of the prompt module within the backbone network. Red arrows denote the insertion locations of of the prompt module.
@@
Fig. 5.: Two-stage training strategy.
3.4 Loss function
@@
where $L_{PR}$ is the triplet loss, used to supervise the place recognition task. $L_{KD}$ is the knowledge distillation loss, designed to ensure that the outputs of replay samples remain consistent between the updated model and the historical model. $\lambda^{\gamma}$ is the weight of $L_{KD}$, $\gamma$ is epoch index, and $\lambda^{\gamma}$ monotonically decreases with $\gamma$. For more details, please refer to InCloud (Knights et al., 2022).
4 Experiment
4.1 Datasets and settings
Datasets. The experimental data encompasses large-scale public datasets (Oxford RobotCar (Maddern et al., 2017), MulRan (Kim et al., 2020), and In-house (Uy and Lee, 2018)) and our self-collected large-scale heterogeneous point cloud dataset (refer to PatchAugNet (Zou et al., 2023)). Specifically, MulRan consists of two parts: DCC and Riverside, while the heterogeneous point cloud dataset comprises two sections: Hankou and WHU-Campus. These datasets were collected in different cities using various mounting platforms and LiDAR sensors. All experimental data have undergone ground removal and coordinate normalization, with each point cloud submap containing 4,096 points. We incrementally train the PCPR models on two dataset sequences: Seq1 (Oxford-¿DCC-¿Riverside-¿In-house) and Seq2 (Oxford-¿Hankou-¿WHU-Campus-¿In-house) respectively. Compared with Seq1, Seq2 exhibits greater data distribution differences, i.e. larger domain gap. More details are shown in Table 1.
Table 1: Details of the experimental data.
@@ 4.1 Datasets and settings
WHU-Campus Campus road (Wuhan)
@@
Baselines. We employ three representative methods for PCPR as backbone networks, namely PointNetVLAD (Uy and Lee, 2018), PatchAugNet [^2] For convenience, the patch feature augmentation module is not used. (Zou et al., 2023), and Minkloc3D (Komorowski, 2021). For CL baselines, we compare against Fine-tuning (denoted as FT), InCloud (Knights et al., 2022), and CCL (Cui and Chen, 2023). Among these, InCloud is the first CL method tailored for PCPR, mitigating catastrophic forgetting through sample replay and knowledge distillation. CCL represents the current SOTA (state-of-the-art) method, it mitigates catastrophic forgetting by introducing contrastive learning to enhance the network’s generalization ability.
Evaluation metrics. Referring to InCloud (Knights et al., 2022), we use the mean top-1 recall ($mR@1_{k}$) and forgetting score ($F$) to quantitatively evaluate the performance of each method in mitigating catastrophic forgetting. Here, $mR@1_{k}$ represents the average top-1 recall of the model on the seen datasets after completing the learning of the $k$th stage. $F$ is a quantitative assessment of the degree of forgetting of history knowledge after finishing training on all datasets. A lower value is preferable. Additionally, we introduce a new metric, i.e., mean incremental top-1 recall ($mIR@1_{k}$), which denotes the average top-1 recall in seen scenes during continual learning (Wang et al., 2024). Compared to $mR@1_{k}$, this metric comprehensively considers the performance of each learning stage. In the CL task for PCPR, after completing the $t$th learning stage, a new model $\mathrm{F}_{t}(x)$ and $\mathrm{P}_{t}(x)$ are obtained. The corresponding top-1 recall $Recall_{t,1:t}$ are acquired on the first $t$ test sets. Upon completing training on all $T$ datasets, a recall matrix $Recall$ is obtained. The specific definitions of three metrics are as follows:
@@
4.2 Quantitative evaluation
Quantitative evaluation on two sequences. Table 2 and Fig. 4.2 are the quantitative evaluation results of all methods on Seq1. As presented in Table 2 and Fig. 4.2, LifelongPR consistently achieves the best performance when incrementally trained on dataset sequence Seq1 with different backbone networks. Compared with the SOTA method CCL, LifelongPR demonstrates significant improvements, with a 3.43% improvement in $mIR@1$, a 6.51% improvement in $mR@1$, and a 2.04% reduction in $F$. Notably, LifelongPR demonstrates particularly substantial performance improvement when employing PointNetVLAD as the backbone network. When employing Minkloc3D as the backbone network, LifelongPR achieves best CL results in terms of recall, surpassing CCL by 1.82% in $mIR@1$, 0.57% in $mR@1$. Although LifelongPR is 1.58% lower than CCL on the $F$ metric, this occurs because CCL overfits on Oxford, degrading its initial performance on DCC and consequently leading to a smaller F.
Table 2: Continual learning results on Seq1. The red and green numbers indicate the performance gains of our method over the SOTA approach (CCL).
Backbone CL Method mIR@1(%$\uparrow$ mR@1(%$\uparrow$ F(%$\downarrow$
@@ 4.2 Quantitative evaluation
Fig. 6.: Average Recall@1 on Seq1. Columns 1-3 correspond to different backbone networks (PointNetVLAD / PatchAugNet / Minkloc3D), and rows 1-4 correspond to different continual learning methods (FT / InCloud / CCL / LifelongPR).
Table 3 and Fig. 4.2 are the quantitative evaluation results of all methods on Seq2. As shown in Table 3 and Fig. 4.2, LifelongPR continues to achieve SOTA performance on the more challenging dataset sequence Seq2. Compared with the SOTA method CCL, it demonstrates significant improvements of 6.50% in $mIR@1$, 7.96% in $mR@1$, while reducing $F$ by 8.95%. Notably, the performance improvements are most pronounced when employing Minkloc3D as the backbone network. With Minkloc3D, LifelongPR attains best CL results, despite not employing contrastive learning like CCL to enhance network generalization and mitigate catastrophic forgetting.
Table 3: Continual learning trained on Seq2. The red and green numbers indicate the performance gains of our method over the SOTA approach (CCL).
Backbone CL Method mIR@1(%$\uparrow$ mR@1(%$\uparrow$ F(%$\downarrow$
@@
Fig. 7.: Average Recall@1 on Seq2. Columns 1-3 correspond to different backbone networks (PointNetVLAD / PatchAugNet / Minkloc3D), and rows 1-4 correspond to different continual learning methods (FT / InCloud / CCL / LifelongPR).
@@
t-SNE visualization. Fig. 4.2 is the t-SNE visualization of samples in Seq2 at different stages of CL. As shown in row 1, after training on Oxford, samples of Oxford exhibit a relatively dispersed distribution in feature space, indicating strong discriminability. However, after finishing CL on Seq2, samples of Oxford gradually become more clustered in feature space, showing reduced discriminability. This demonstrates catastrophic forgetting. Rows 2-4 show that sample replay, knowledge distillation, and prompt learning effectively mitigate catastrophic forgetting of the PCPR model. Among these methods, CCL and the proposed method LifelongPR perform the best, maintaining dispersed distributions of samples from each dataset in feature space across different learning stages. It demonstrates that our method has strong anti-forgetting capabilities, consistent with quantitative results in Table 3 and Fig. 4.2.
Fig. 8.: t-SNE visualization of samples in Seq2 at different stages of continuous learning. Rows 1-4 corresponds to different continual learning methods (FT / InCloud / CCL / LifelongPR).
Success cases. Fig. 4.2 presents six success cases of our method in typical street scenes. In cases 1-4, the model trained on Oxford successfully retrieves the correct top-1 submaps. However, retrieval fails after CL using FT on Seq1 and Seq2, whereas it succeeds when LifelongPR is used. In cases 5-6, the model trained on Oxford fails to retrieve the correct top-1 submaps, and retrieval still fails after CL using FT on Seq1 and Seq2, but succeeds when LifelongPR is used. Analysis of specific cases reveals that the original PCPR model has strong capabilities in extracting scene detail information. After CL using FT, the model’s local feature extraction ability significantly declines, leading to retrieval failures. In contrast, our method effectively mitigates the decline in feature extraction capability through more reasonable replay sample selection and a prompt learning-based CL framework.
Fig. 9.: Six success cases of LifelongPR. Column 1: query; Column 2: top 1 submaps retrieved by the model trained on Oxford; Columns 3-4: top 1 submaps retrieved by the model trained on Seq1; Columns 5-6: top 1 submaps retrieved by the model trained on Seq2.
5 Analysis and discussion
@@
5.1 Effectiveness of replay sample selection method
Table 4 presents the ablation experiment results of the core modules. (a) utilizes random sampling for replay sample selection, yielding poor CL performance. In contrast, (b), (c), and (d) demonstrate significant performance improvements, with $mIR@1$ increasing by 0.31%–0.69% and $mR@1$ by 0.95%–1.36%, while $F$ decreases by 1.38%-4.11%. This performance improvement stems from our sophisticated replay sample selection method, which considers sample distributions in both Euclidean and feature spaces, thereby identifying and retaining more informative samples for replaying. On the basis of (d), (e) further enhances performance by incorporating information quantity differences across datasets when allocating replay sample size, achieving additional improvements of 0.89% in $mIR@1$ and 0.60% in $mR@1$. Fig. 5.1 (b) shows the replay sample size allocation results. Compared to uniformly allocating replay sample sizes across all training sets, our method allocates a greater number of replay samples to more informative training sets, such as Oxford and In-house, enabling more effective utilization of limited replay sample budget.
Table 4: Ablation experiment results of key modules including replay sample selection and the prompt module. Random: randomly sampling; Greedy: sampling based on greedy algorithm; InfoQ: allocating sample size based on information quantity.
ID Select Forget Prompt mIR@1(%$\uparrow$ mR@1(%$\uparrow$ F(%$\downarrow$
@@ 5.1 Effectiveness of replay sample selection method
Fig. 10.: Replay sample size allocation results.
5.2 Effectiveness of prompt module
As shown in Table 4, on the basis of (e), (f) further enhances performance by employing the prompt module to extract domain-specific knowledge from each dataset and guiding the backbone network to to extract features adapted to samples. This yields additional improvements of 1.13% in $mIR@1$ and 2.28% in $mR@1$, while reducing $F$ by 4.17%. During CL, the prompt module effectively serves as an external memory unit. By combining a two-stage training strategy with replay samples, it successfully captures distinctive domain-specific knowledge from each dataset. At inference time, the encoded representations generated by the prompt module are fed as auxiliary inputs to the backbone network, directing it to extract sample-adaptive global features.
5.3 Effectiveness of two-stage training strategy
Table 5 and Fig. 5.3 present CL results under different training strategies, where our two-stage strategy achieves best performance while training only the prompt module performs worst. Comparative analysis of Fig. 5.3 (a)-(c) demonstrates that the naive addition of a prompt module fails to improve the model’s anti-forgetting capabilities, regardless of one-stage joint training or prompt-only training. As shown in Table 5 and Fig. 5.3 (d), the two-stage training strategy shows significant improvements over the one-stage training strategy, with gains of 3.17% in $mIR@1$, 6.22% in $mR@1$, and a 7.83% reduction in $F$. In stage 1, the prompt module is trained to extract domain-specific knowledge from the current dataset, while in stage 2, the backbone network is fine-tuned to adapt to the target domain of the training set. The experimental results demonstrate the rationality and effectiveness of this two-stage training strategy.
Table 5: Continual learning results of LifelongPR with different training strategies.
Train Strategy mIR@1(%$\uparrow$ mR@1(%$\uparrow$ F(%$\downarrow$

### v2

@@ III-B1 Sample size allocation based on information quantity
[…]
where $\gamma$ is a parameter of the Gaussian kernel function, and $f_{i}$ is the feature of the $i$th sample in $D$. Then, we use singular value decomposition (SVD) to estimate the effective rank of $A$ as follows:
@@
III-B2 Sample selection based on spatial distribution
@@
[…]
@@ 22 $M_{1:t}=M_{1:t-1}\cup\{M_{t}\}$;
III-C CL framework based on prompt learning
@@
Prompt module. We design the prompt module with reference to QFormer [56]. It consists of two MLPs (denoted as $\mathrm{MLP_{in}}:R^{3}\rightarrow R^{d}$ and $\mathrm{MLP_{out}}:R^{d}\rightarrow R^{d^{\prime}}$), $k_{q}$ learnable prompts (denoted as $Q\in R^{k_{q}\times d}$), and $N$ attention modules (denoted as $\mathrm{Attn}:R^{d}\rightarrow R^{d}$), as shown in Fig. 2 (c). Here, $\mathrm{MLP_{in}}$ maps the raw point cloud (denoted as $x\in R^{N_{x}\times 3}$) into a high-dimensional space, which then serves as the input to the attention modules. The learnable prompts $Q$ store domain-specific information from each training set and act as the Query in the attention modules. The attention modules $\mathrm{Attn}$ facilitate interaction between $\mathrm{MLP_{in}}(x)$ and $Q$ to extract domain-specific information related to $x$. $\mathrm{MLP_{out}}$ adjusts the dimensionality of the attention module outputs, making it easier to embed prompt information into the backbone network. The dimension $d^{\prime}$ is tailored to accommodate different backbone networks. The extraction of domain-specific information from raw point cloud $x$ by the prompt module is formally expressed as:
@@
Determining how to integrate learnable prompts into the backbone network is a critical problem. Existing prompt learning methods are primarily designed for transformer-based networks. However, most deep learning-based PCPR methods rely on sparse convolution or PointNet-style architectures, requiring the insertion location of prompts to be tailored to the specific structure of the backbone network. For Minkloc3D, a PCPR method based on sparse convolution, this study directly serializes $\mathrm{P}(x)$ with the raw point cloud and uses it as the network input. For PCPR models based on PointNet, such as PointNetVLAD and PatchAugNet, $\mathrm{P}(x)$ is directly added to the intermediate features of the backbone network. Crucially, our method requires no modifications to backbone networks. The insertion locations of the prompt module are visualized in Fig. 4.
Figure 4: Location of the prompt module within the backbone network. (a) PointNetVLAD, (b) Minkloc3D, (c) PatchAugNet. Red arrows denote the insertion locations of of the prompt module.
@@
Figure 5: Two-stage training strategy. (a) stage 1, (b) stage 2.
III-D Loss function
@@
where $L_{PR}$ is the triplet loss, used to supervise the place recognition task. $L_{KD}$ is the knowledge distillation loss, designed to ensure that the outputs of replay samples remain consistent between the updated model and the historical model. $\lambda^{\gamma}$ is the weight of $L_{KD}$, $\gamma$ is epoch index, and $\lambda^{\gamma}$ monotonically decreases with $\gamma$. For more details, please refer to InCloud [19].
IV Experiment
IV-A Datasets and settings
[…]
TABLE I: Details of the experimental data.
@@ IV-A Datasets and settings
Campus Campus road (Wuhan)
@@
Baselines. We employ three representative methods for PCPR as backbone networks, namely PointNetVLAD [11], PatchAugNet [^2] For convenience, the patch feature augmentation module is not used. [13], and Minkloc3D [12]. For CL baselines, we compare against Fine-tuning (denoted as FT), InCloud [19], and CCL [20]. Among these, InCloud is the first CL method tailored for PCPR, mitigating catastrophic forgetting through sample replay and knowledge distillation. CCL represents the current SOTA (state-of-the-art) method, it mitigates catastrophic forgetting by introducing contrastive learning to enhance the network’s generalization ability.
Evaluation metrics. Referring to InCloud [19], we use the mean top-1 recall ($mR@1_{k}$) and forgetting score ($F$) to quantitatively evaluate the performance of each method in mitigating catastrophic forgetting. Here, $mR@1_{k}$ represents the average top-1 recall of the model on the seen datasets after completing the learning of the $k$th stage. $F$ is a quantitative assessment of the degree of forgetting of history knowledge after finishing training on all datasets. A lower value is preferable. Additionally, we introduce a new metric, i.e., mean incremental top-1 recall ($mIR@1_{k}$), which denotes the average top-1 recall in seen scenes during continual learning [35]. Compared to $mR@1_{k}$, this metric comprehensively considers the performance of each learning stage. In the CL task for PCPR, after completing the $t$th learning stage, a new model $\mathrm{F}_{t}(x)$ and $\mathrm{P}_{t}(x)$ are obtained. The corresponding top-1 recall $Recall_{t,1:t}$ are acquired on the first $t$ test sets. Upon completing training on all $T$ datasets, a recall matrix $Recall$ is obtained. The specific definitions of three metrics are as follows:
@@
IV-B Quantitative evaluation
Quantitative evaluation on two sequences. Table II and Fig. 6 are the quantitative evaluation results of all methods on Seq1. As presented in Table II and Fig. 6, LifelongPR consistently achieves the best performance when incrementally trained on dataset sequence Seq1 with different backbone networks. Compared with the SOTA method CCL, LifelongPR demonstrates significant improvements, with a 3.43% improvement in $mIR@1$, a 6.51% improvement in $mR@1$, and a 2.04% reduction in $F$. Notably, LifelongPR demonstrates particularly substantial performance improvement when employing PointNetVLAD as the backbone network. When employing Minkloc3D as the backbone network, LifelongPR achieves best CL results in terms of recall, surpassing CCL by 1.82% in $mIR@1$, 0.57% in $mR@1$. Although LifelongPR is 1.58% lower than CCL on the $F$ metric, this occurs because CCL overfits on Oxford, degrading its initial performance on DCC and consequently leading to a smaller F.
TABLE II: Continual learning results on Seq1. The red and green numbers indicate the performance gains of our method over the SOTA approach (CCL).
[…]
@@ IV-B Quantitative evaluation
Figure 6: Average Recall@1 on Seq1. (a) FT, (b) InCloud, (c) CCL (SOTA method), (d) LifelongPR (ours), (d) improvements of LifelongPR compared to CCL. Rows 1-3 correspond to different backbone networks.
Table III and Fig. 7 are the quantitative evaluation results of all methods on Seq2. As shown in Table III and Fig. 7, LifelongPR continues to achieve SOTA performance on the more challenging dataset sequence Seq2. Compared with the SOTA method CCL, it demonstrates significant improvements of 6.50% in $mIR@1$, 7.96% in $mR@1$, while reducing $F$ by 8.95%. Notably, the performance improvements are most pronounced when employing Minkloc3D as the backbone network. With Minkloc3D, LifelongPR attains best CL results, despite not employing contrastive learning like CCL to enhance network generalization and mitigate catastrophic forgetting.
TABLE III: Continual learning trained on Seq2. The red and green numbers indicate the performance gains of our method over the SOTA approach (CCL).
[…]
@@
Figure 7: Average Recall@1 on Seq2. (a) FT, (b) InCloud, (c) CCL (SOTA method), (d) LifelongPR (ours), (d) improvements of LifelongPR compared to CCL. Rows 1-3 correspond to different backbone networks.
Fig. 8 presents the place recognition results on Oxford after continual learning on Seq1. As shown in Fig. 8 (a), after continual learning on Seq1 with FT, top-1 recall drop from 95.10% to 72.00%, declining by 23.10%. As shown in Fig. 8 (b), after continual learning on Seq1 with CCL, top-1 recall drop from 94.10% to 83.20%, declining by 10.90%. As shown in Fig. 8 (c), after continual learning on Seq1 with LifelongPR, top-1 recall drop from 95.10% to 87.90%, declining by only 7.20%. Our method LifelongPR achieves the best performance, exhibiting excellent anti-forgetting capabilities.
Figure 8: Place recognition results on Oxford after continual learning on Seq1. (a) FT ($\downarrow$23.10%), (b) CCL ($\downarrow$10.90%), (d) LifelongPR (ours, $\downarrow$7.20%). $R@1$ means top-1 recall and only 1000 queries are selected from Oxford for better representation.
@@
t-SNE visualization. Fig. 9 is the t-SNE visualization of samples in Seq2 at different stages of CL. As shown in row 1, after training on Oxford, samples of Oxford exhibit a relatively dispersed distribution in feature space, indicating strong discriminability. However, after finishing CL on Seq2, samples of Oxford gradually become more clustered in feature space, showing reduced discriminability. This demonstrates catastrophic forgetting. Rows 2-4 show that sample replay, knowledge distillation, and prompt learning effectively mitigate catastrophic forgetting of the PCPR model. Among these methods, CCL and the proposed method LifelongPR perform the best, maintaining dispersed distributions of samples from each dataset in feature space across different learning stages. It demonstrates that our method has strong anti-forgetting capabilities, consistent with quantitative results in Table III and Fig. 7.
Figure 9: t-SNE visualization of samples in Seq2 at different stages of continuous learning. Rows 1-4 corresponds to different continual learning methods, (a) FT, (b) InCloud, (b) CCL, (d) LifelongPR (ours).
Success cases. Fig. 10 presents three success cases of our method in typical street scenes. In cases 1-2, the model trained on Oxford successfully retrieves the correct top-1 submaps. However, retrieval fails after CL using FT, whereas it succeeds when LifelongPR is used. In case 3, the model trained on Oxford fails to retrieve the correct top-1 submap, and retrieval still fails after CL using FT, but succeeds when LifelongPR is used. Analysis of specific cases reveals that the original PCPR model has strong capabilities in extracting scene detail information. After CL using FT, the model’s local feature extraction ability significantly declines, leading to retrieval failures. In contrast, our method effectively mitigates the decline in feature extraction capability through more reasonable replay sample selection and a prompt learning-based CL framework.
Figure 10: Three success cases of LifelongPR on Seq2. Column 1: query; Column 2: top 1 submaps retrieved by the model trained on Oxford; Columns 3-4: top 1 submaps retrieved by the model trained on Seq2.
V Analysis and discussion
@@
V-A Effectiveness of replay sample selection method
Table IV presents the ablation experiment results of the core modules. (a) utilizes random sampling for replay sample selection, yielding poor CL performance. In contrast, (b), (c), and (d) demonstrate significant performance improvements, with $mIR@1$ increasing by 0.31%–0.69% and $mR@1$ by 0.95%–1.36%, while $F$ decreases by 1.38%-4.11%. This performance improvement stems from our sophisticated replay sample selection method, which considers sample distributions in both Euclidean and feature spaces, thereby identifying and retaining more informative samples for replaying. On the basis of (d), (e) further enhances performance by incorporating information quantity differences across datasets when allocating replay sample size, achieving additional improvements of 0.89% in $mIR@1$ and 0.60% in $mR@1$. Fig. 11 (b) shows the replay sample size allocation results. Compared to uniformly allocating replay sample sizes across all training sets, our method allocates a greater number of replay samples to more informative training sets, such as Oxford and In-house, enabling more effective utilization of limited replay sample budget.
TABLE IV: Ablation experiment results of key modules including replay sample selection and the prompt module. Random: randomly sampling; Greedy: sampling based on greedy algorithm; InfoQ: allocating sample size based on information quantity.
[…]
@@ V-A Effectiveness of replay sample selection method
Figure 11: Replay sample size allocation results. (a) not use information quantity, (b) use information quantity (ours).
V-B Effectiveness of prompt module
As shown in Table IV, on the basis of (e), (f) further enhances performance by employing the prompt module to extract domain-specific knowledge from each dataset and guiding the backbone network to to extract features adapted to samples. This yields additional improvements of 1.13% in $mIR@1$ and 2.28% in $mR@1$, while reducing $F$ by 4.17%. During CL, the prompt module effectively serves as an external memory unit. By combining a two-stage training strategy with replay samples, it successfully captures distinctive domain-specific knowledge from each dataset. At inference time, the encoded representations generated by the prompt module are fed as auxiliary inputs to the backbone network, directing it to extract sample-adaptive global features.
V-C Effectiveness of two-stage training strategy
Table V and Fig. 12 present CL results under different training strategies, where our two-stage strategy achieves best performance while training only the prompt module performs worst. Comparative analysis of Fig. 12 (a)-(c) demonstrates that the naive addition of a prompt module fails to improve the model’s anti-forgetting capabilities, regardless of one-stage joint training or prompt-only training. As shown in Table V and Fig. 12 (d), the two-stage training strategy shows significant improvements over the one-stage training strategy, with gains of 3.17% in $mIR@1$, 6.22% in $mR@1$, and a 7.83% reduction in $F$. In stage 1, the prompt module is trained to extract domain-specific knowledge from the current dataset, while in stage 2, the backbone network is fine-tuned to adapt to the target domain of the training set. The experimental results demonstrate the rationality and effectiveness of this two-stage training strategy.
TABLE V: Continual learning results of LifelongPR with different training strategies.
[…]